# PICore: Physics-Informed Unsupervised Coreset Selection for Data Efficient Neural Operator Training

### v1

@@ C.1 Zero Shot Super Resolution
graNd $1.77\times 10^{0}$ $1.57\times 10^{0}$ $1.59\times 10^{0}$ $1.55\times 10^{0}$ $1.52\times 10^{0}$ $1.72\times 10^{0}$
@@
gradmatch $1.70\times 10^{-1}$ $7.83\times 10^{-2}$ $7.36\times 10^{-2}$ $7.61\times 10^{-2}$ $7.76\times 10^{-2}$ $8.54\times 10^{-2}$
@@
graNd $1.55\times 10^{-1}$ $7.83\times 10^{-2}$ $7.57\times 10^{-2}$ $7.53\times 10^{-2}$ $7.75\times 10^{-2}$ $8.54\times 10^{-2}$
UNO Supervised craig $5.99\times 10^{-2}$ $5.79\times 10^{-2}$ $5.64\times 10^{-2}$ $5.56\times 10^{-2}$ $5.52\times 10^{-2}$ $5.16\times 10^{-2}$
gradmatch $5.99\times 10^{-2}$ $5.83\times 10^{-2}$ $5.68\times 10^{-2}$ $5.60\times 10^{-2}$ $5.53\times 10^{-2}$ $5.16\times 10^{-2}$
adacore $6.45\times 10^{-2}$ $5.98\times 10^{-2}$ $5.77\times 10^{-2}$ $5.58\times 10^{-2}$ $5.49\times 10^{-2}$ $5.16\times 10^{-2}$
el2n $5.93\times 10^{-2}$ $5.77\times 10^{-2}$ $5.66\times 10^{-2}$ $5.57\times 10^{-2}$ $5.52\times 10^{-2}$ $5.16\times 10^{-2}$
graNd $5.94\times 10^{-2}$ $5.82\times 10^{-2}$ $5.69\times 10^{-2}$ $5.60\times 10^{-2}$ $5.52\times 10^{-2}$ $5.16\times 10^{-2}$
PICore craig $6.07\times 10^{-2}$ $5.74\times 10^{-2}$ $5.66\times 10^{-2}$ $5.53\times 10^{-2}$ $5.51\times 10^{-2}$ $5.16\times 10^{-2}$
gradmatch $5.97\times 10^{-2}$ $5.76\times 10^{-2}$ $5.67\times 10^{-2}$ $5.60\times 10^{-2}$ $5.55\times 10^{-2}$ $5.16\times 10^{-2}$
adacore $6.56\times 10^{-2}$ $6.02\times 10^{-2}$ $5.81\times 10^{-2}$ $5.57\times 10^{-2}$ $5.50\times 10^{-2}$ $5.16\times 10^{-2}$
el2n $5.96\times 10^{-2}$ $5.76\times 10^{-2}$ $5.68\times 10^{-2}$ $5.58\times 10^{-2}$ $5.53\times 10^{-2}$ $5.16\times 10^{-2}$
graNd $5.96\times 10^{-2}$ $5.77\times 10^{-2}$ $5.65\times 10^{-2}$ $5.58\times 10^{-2}$ $5.54\times 10^{-2}$ $5.16\times 10^{-2}$

### v2

@@ C.5 Zero Shot Super Resolution
graNd $1.77\times 10^{0}$ $1.57\times 10^{0}$ $9\times 10^{0} 1.55×1$$5\times 10^{0}$ $1.52\times 10^{0}$ $1.72\times 10^{0}$
@@
gradmatch $1.70\times 10^{-1}$ $7.83\times 10^{-2}$ $7.36\times 10^{-2}$ $61\times 10^{-2} 7.7$$76\times 10^{-2}$ $8.54\times 10^{-2}$
@@
graNd $1.55\times 10^{-1}$ $7.83\times 10^{-2}$ $7.57\times 10^{-2}$ $53\times 10^{-2} 7.7$$75\times 10^{-2}$ $8.54\times 10^{-2}$
UNO Supervised craig $5.99\times 10^{-2}$ $5.79\times 10^{-2}$ $64\times 10^{-2} 5.5$$6\times 10^{-2} 5.5$$52\times 10^{-2}$ $5.16\times 10^{-2}$
gradmatch $5.99\times 10^{-2}$ $5.83\times 10^{-2}$ $5.68\times 10^{-2}$ $60\times 10^{-2} 5.5$$53\times 10^{-2}$ $5.16\times 10^{-2}$
adacore $6.45\times 10^{-2}$ $5.98\times 10^{-2}$ $77\times 10^{-2} 5.5$$58\times 10^{-2}$ $5.49\times 10^{-2}$ $5.16\times 10^{-2}$
el2n $5.93\times 10^{-2}$ $5.77\times 10^{-2}$ $66\times 10^{-2} 5.5$$7\times 10^{-2} 5.5$$52\times 10^{-2}$ $5.16\times 10^{-2}$
graNd $5.94\times 10^{-2}$ $5.82\times 10^{-2}$ $5.69\times 10^{-2}$ $60\times 10^{-2} 5.5$$52\times 10^{-2}$ $5.16\times 10^{-2}$
PICore craig $6.07\times 10^{-2}$ $5.74\times 10^{-2}$ $66\times 10^{-2} 5.5$$3\times 10^{-2} 5.5$$51\times 10^{-2}$ $5.16\times 10^{-2}$
gradmatch $5.97\times 10^{-2}$ $5.76\times 10^{-2}$ $5.67\times 10^{-2}$ $60\times 10^{-2} 5.5$$55\times 10^{-2}$ $5.16\times 10^{-2}$
adacore $6.56\times 10^{-2}$ $6.02\times 10^{-2}$ $81\times 10^{-2} 5.5$$7\times 10^{-2} 5.5$$50\times 10^{-2}$ $5.16\times 10^{-2}$
el2n $5.96\times 10^{-2}$ $5.76\times 10^{-2}$ $68\times 10^{-2} 5.5$$8\times 10^{-2} 5.5$$53\times 10^{-2}$ $5.16\times 10^{-2}$
graNd $5.96\times 10^{-2}$ $5.77\times 10^{-2}$ $65\times 10^{-2} 5.5$$8\times 10^{-2} 5.5$$54\times 10^{-2}$ $5.16\times 10^{-2}$